OpenStudy (anonymous):

If f(x) = 2x + 1 and g(x) = −2x − 3, find (f + g)(−1).

OpenStudy (anonymous):

(f + g)(-1) = f(-1) + g(-1)

OpenStudy (anonymous):

so i enter =1 into both equations?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

f(-1)+g(-1)=-1-1 = -2

OpenStudy (anonymous):

okay thanks! I think i got it!